hi can you help me solve this twp step equation: 2+k/5 =10
Answer by mananth(16946)

2 +k/5 =10
add -2
2-2 +k/5 =10-2
k/5 = 8
multiply my by 5
k/5 * 5 = 8*5
k=40
...
m.ananth@hotmail.ca